A kidnapper demands a 44.0-kg cube of gold as a ransom. The density of gold is 19.3×103kg/m3.
Answer #1

muss density = volume

19.3\times 10^{3} = \frac{44}{volume}

volume = \frac{44}{19.3\times 10^{3}} = 2.279\times 10^{-3} m^{3}

volume = a^{3}

a^{3} = 2.279\times 10^{-3}

a = 0.1332m